Solve: logb 36=-2. b is the small number next to log.
Answer by scott8148(6628)

using rules of logarithms __ b^(-2)=36

using rules of exponents __ 1/(b^2)=36

multiplying by b^2 and dividing by 36 __ 1/36=b^2

taking sqrt __ 1/6=b